# Network-Level Prompt and Trait Leakage in Local Research Agents

### v2

@@ 3.1 Threat Model
Adversary Goals. The adversary’s objective is to infer sensitive information about the user from the browsing trace. We focus on two representative leakage goals that capture both short- and long-term risks. The first is prompt recovery, where the adversary seeks to infer queries functionally equivalent to the user’s original input (e.g., detecting that a user asked about “signs of depression and seeking professional mental health support”). The second is trait inference, where the adversary profiles latent attributes such as gender, religion, and political ideology by correlating patterns across multiple browsing sessions. Together, these goals illustrate how metadata can expose both immediate user intent and persistent personal characteristics.
Adversary Capabilities. The adversary can observe domain names accessed by the agent, obtained through plaintext DNS or TLS handshake metadata such as Server Name Indication (SNI). In addition, the adversary can track the sequence of domains accessed, capturing the order of domain visits during the agent’s browsing sessions, as well as packet-level metadata including timing and payload lengths. By contrast, the adversary does not see full URLs, such as query strings (e.g., /search?q=...), page contents including articles or forms, the plaintext of the user’s original prompts, or the agent’s intermediate queries. Visibility may also be reduced by Content Delivery Networks (CDNs), reverse proxies, or load balancers that aggregate or mask requests. To capture these cases, we also consider partial-visibility settings where only a subset of domains (e.g., 80%, 60%) is observable. These capabilities align with realistic adversaries such as local ISPs, corporate firewalls, VPN providers, or other on-path observers that collect encrypted metadata without decryption or manipulation.
Adversary Assumptions. We assume the user interacts with a local web or research agent that autonomously performs multi-hop web browsing in response to a natural language prompt. The adversary is aware of the agent’s general browsing capabilities, such as issuing web searches, clicking links, scrolling pages, or scraping structured data from web pages. We further assume that the adversary observes the entire session between the agent and the internet rather than isolated flows, that DNS and TLS handshake metadata remain visible (i.e., no DoH, DoT, or ECH), and that users do not employ VPNs, Tor, or other anonymizing proxies that would conceal domain sequences completely. Even under full-session monitoring, however, proxy infrastructures such as CDNs may still limit visibility, exposing only a subset of traces.
@@ 3.2 Attack Scenarios
Academic WRAs. In university, students and faculty use local WRAs for daily tasks, coursework, or research, leaking information about health, politics, or unpublished work. External observers such as ISPs or competing research entities could exploit traces to reconstruct prompts or profile academic interests.
@@ 4 Attack Methodology
We present the end-to-end pipeline for our privacy attacks against AI agents. Both tasks—prompt recovery and trait inference—follow a shared two-stage process: (1) offline model construction on proxy datasets or , yielding or ; and (2) online attack execution on real user traces or to produce or . This separation reflects adversary constraints: real user prompts are private and unavailable for training, so the adversary must first build an inference model on proxy data before deploying it. The pipeline remains structurally consistent across tasks, though the input sources and inference targets differ.
@@ 4.1 Proxy Prompt Generation
To train the inference model in Stage 1, the adversary generates proxy prompts that realistically simulate user behavior and produce informative traces . Effective proxy prompts must balance realism, browsing induction, and inferential utility. (a) Prompts resemble natural-language queries real users might issue (health, education, planning), forming or . (b) They must reliably trigger multi-page browsing, rather than being satisfied by a single source (e.g., Wikipedia) or by the agent’s prior knowledge; to enforce this, we prepend instructions requiring at least five distinct page visits. (c) Prompts must be sufficiently detailed to produce distinguishable traces; underspecified inputs are rewritten into instruction-complete forms to avoid trivial queries that yield little signal. (d) Prompts and resulting traces remain semantically linked, so the trace encodes sufficient signal for recovering or inferring .
Prompt recovery. We draw proxy prompts from public query datasets (FEDWEB13, SESSION14, DD16) to build . Some prompts are underspecified to reliably induce meaningful browsing, and agents differ in how they handle such inputs. GPT-Researcher [21] internally refines vague prompts, but ChatGPT Deep Research UI asks clarifying questions, while Deep Research API, AutoGen, and Browser-Use execute inputs as-is. To align behaviors, Deep Research API prompts are rewritten into detailed variants (denoted -DR) to match the interactive Deep Research interface.
Trait inference. From persona profiles, we construct , embedding selected traits (e.g., age, citizenship, insurance coverage) either explicitly or implicitly into prompts. This mirrors real user behavior, where personal context guides information seeking (e.g., health questions tied to age or history). A persistent adversary linking sessions over time (e.g., by IP) would observe recurring patterns tied to these traits (). Executing such prompts through agents produces labeled traces that reveal stable user attributes, enabling training of trait inference models (). Dataset-specific sampling and generation details are described in Section 6.1.
@@ 4.2 Domain Trace Collection
To understand how different prompts yield distinct browsing behavior, we collect domain-level metadata from agents via Playwright logs or built-in system logs, yielding a trace for each prompt , where capture the ordered sequence of accessed domains, timestamps (at millisecond-level precision), and payload size in bytes, corresponding to the life cycle of a single user prompt. This captures whether the agent revisits the same domains or explores diverse sources, the features that adversaries could exploit. Although TLS application data does not expose full URLs or domain names directly, domain attribution remains feasible through DNS queries, SNI fields in TLS handshakes, or correlated DNS/HTTPS timing, all of which are observable to ISPs or other on-path entities. This matches realistic adversarial capabilities today, though we note that CDN consolidation or future deployment of DoH/DoT/ECH could reduce visibility.
Noise Filtering. A passive adversary observes all outbound requests, including primary domains, embedded resources, ads, and trackers. Because much of this traffic is incidental rather than agent-driven, an effective adversary would attempt to filter noise to isolate the domains most indicative of user intent. We apply two simple heuristics for the noise filtering: removing well-known advertising and analytics services (e.g., doubleclick.net, onetrust.com, taboola.com), and discarding packets smaller than 7KB, which in pilot runs rarely corresponded to substantive browsing activity. These steps represent the simplest plausible filtering approach; more advanced adversaries could employ refined techniques (e.g., traffic classification or resource-type analysis) to separate incidental requests from meaningful domains. We leave such stronger filtering strategies to future work.
@@ 4.3 Inference Model Construction and Use
For prompt recovery, the task is to reconstruct a natural-language prompt that captures the functional intent of the user’s original input. The adversary constructs an inference model () from proxy (trace, prompt) pairs in the proxy dataset (), using either ICL, where a small set of examples is included in the inference prompt without parameter updates, or fine-tuning, where the same proxy pairs are used to train a direct mapping from traces to prompts. Given a new user prompt producing trace (optionally filtered to per Section 4.2), the resulting model is then applied to new traces to reconstruct real user prompts .
@@
5 OBELS: Ontology-aware Behavioral Leakage Score
From the adversary’s perspective, a reconstructed prompt poses a significant privacy risk if it induces the same browsing behavior as the user’s original input. Even when the textual form differs, the sensitive intent is revealed through replicated domain visits and task flows. Existing metrics focus on textual resemblance and therefore miss such cases. We introduce OBELS (Ontology-aware Behavioral Leakage Scores), a multi-metric scoring scheme that decomposes leakage into dimensions of preserved functionality, domain targets, and semantic entities, providing a structured alignment with attacker goals and real-world agent behavior.
@@ 5.1 Limitations of Existing Metrics
In AI agent contexts, leakage risk stems from functional equivalence: a prompt recovery is risky if it causes the agent to visit similar sites, access comparable services, or pursue equivalent goals—regardless of its phrasing. Existing metrics fall short because they evaluate what prompts say, not what they do.
SBERT score, computed as cosine similarity between sentence embeddings [62], captures high-level semantic alignment but suffers from paraphrastic inflation. It misleadingly assigns high scores to verbose, generic, or loosely related prompts [71, 87], overlooking task-specific behavior or intent.
LLM-as-Judge methods leverage LLMs (e.g., GPT-4) to rate similarity through natural language instructions. While more tolerant to paraphrase and flexible, they are prompt-sensitive and often lack interpretability [61, 36], producing scalar scores without indicating which semantic elements were preserved or lost.
These shortcomings are evident in practice. For example, rephrasing “Compare Italy and Spain’s digital nomad visas” as “How do work visas differ across southern Europe?” yields low cosine similarity, yet the agent navigates to similar government portals and retrieves overlapping content. From a privacy standpoint, this is a successful reconstruction—yet traditional metrics fail to capture it.
@@ 5.2 Semantic Triplet Scoring with OBELS
To address these gaps, we propose OBELS, a set of four complementary scores designed to capture behavioral alignment. Rather than collapsing leakage into a single scalar, OBELS reports distinct dimensions so that failures in one area (e.g., missing entity granularity) are not obscured by high alignment elsewhere. This decomposition provides interpretable diagnostics and better reflects adversarial risk. It does so by abstracting prompts into structured triplets:
@@
Here, intent denotes the high-level goal (e.g., learn, explore, analyze, compare, summarize, plan, decide, book, watch, read, evaluate); source_type identifies the class of information or service sought (e.g., travel, symptom, policy_area, event, visa_process, treatment_method, academic_field, cuisine, etc); and entity specifies the core topic (e.g., Rome, Italy, depression, face transplant, immigration, PhD, Business, Swahili dish, etc). (Figure 10).
@@
Each dimension is scored from 0.0 (unrelated) to 1.0 (fully equivalent) using a standardized evaluation template applied to GPT-4, which compares triplet sets holistically and assigns fine-grained similarity scores (see Figure 11).
Unlike conventional metrics, OBELS captures reconstructions that preserve function and downstream behavior, even when their exact wordings diverge. By decomposing leakage into interpretable components, it offers a more transparent measure of privacy risk in prompt recovery.
@@
| FEDWEB13 [49] | Prompt | 50 prompts | 20 | GPT-Researcher | 77.5 | 143.0 | 1 prompt 1 session |
| SESSION14 [50] | Prompt | 60 prompts | 20 | GPT-Researcher | 69.9 | 145.9 | 1 prompt 1 session |
| DD16 [51] | Prompt | 53 prompts | 20 | GPT-Researcher | 64.8 | 155.8 | 1 prompt 1 session |
| PERSONA [14] | Trait | 50 personas | 3 | AutoGen | 144.1 | 173.7 | 3–5 prompts 7 sessions |
@@ 6.1 Datasets
Prompt Recovery (TREC). We use three TREC datasets: FedWeb 2013 (50 topics) [49], Session 2014 (60) [50], and Dynamic Domain 2016 (53) [51], referred to as FEDWEB13, SESSION14, and DD16. For FEDWEB13 and DD16, we concatenate the <description> and <narrative> fields to match SESSION14’s instruction-style and prompt specificity. To align with OpenAI’s Deep Research API [55], we also generate -DR variants by rewriting each prompt with GPT-4.1, while scoring is always against the original prompts. For each dataset, 20 prompts are reserved for testing, with the remainder used for training and ICL examples.
Trait Inference (PERSONA). We sample 50 profiles from the PERSONA_subset [14] (997 total), each annotated with 32 traits. For each, we select 5 traits and use GPT-4o to generate 3–5 prompts across 7 sessions, simulating a week of browsing. Three personas are held out for ICL, and after filtering to ages 18–70, 35 remain for testing. The prompt template generating queries per user (persona) is provided in Appendix H.
@@ 6.2 WRAs and Backbone LLMs
GPT Researcher [21]. An open-source autonomous research agent for deep multi-hop investigation. We run it in deep mode, which produces structured research traces. GPT Researcher strategically uses multiple LLM backbones, and we evaluate both GPT-family and open-source backbones for the prompt recovery task.
Browser-Use [43]. An open-source web agent that automates browsing through a local browser. To ensure stable traces, we configure it to start from Bing and require at least five distinct page visits before summarization. Browser-Use is used for the prompt recovery task with GPT-4o as the backbone.
AutoGen [84]. An open-source web agent system we used for both prompt recovery (GPT-4o) and trait inference (Gemini 2.0 Flash), instrumented to log domains and metadata. Unlike the other agents, it produces the real packet-level traces, noisier with ads and trackers.
@@ 6.3 Evaluation Metrics
(1) SBERT, computed as cosine similarity over mean-pooled Sentence-BERT (SBERT) embeddings [62];
(2) LLM-as-Judge (template in Figure 12), where an LLM rates prompt similarity based on natural language instructions; and
(3) OBELS, our proposed behavior-aware scoring scheme to capture functional equivalence across multiple dimensions (i.e., whether the reconstructed prompt would induce similar agent behavior).
@@
Categorical traits use exact match for single-token values; multi-word categories are scored using the SBERT score to account for semantic equivalence.
Free-text traits are assessed using the SBERT score, which computes semantic overlap via contextual token embeddings.

### v3

@@ 3.1 Threat Model
Adversary Goals. The adversary’s objective is to infer sensitive user information from the browsing trace. We focus on two representative leakage goals that capture both short- and long-term risks. The first is prompt recovery, where the adversary seeks to infer queries functionally equivalent to the user’s original input (e.g., detecting that a user asked about “signs of depression and seeking professional mental health support”). The second is trait inference, where the adversary profiles latent attributes such as gender, religion, and political ideology by correlating patterns across multiple browsing sessions. Together, these illustrate how metadata can expose both immediate user intent and persistent personal characteristics.
Adversary Capabilities. The adversary observes domain-level metadata generated by the agent’s web activity, including domain names (via plaintext DNS or TLS handshake metadata such as SNI), the order of domain visits, and coarse packet-level features such as timing and payload sizes. The adversary does not observe full URLs (e.g., query strings), page contents, user prompts, or the agent’s intermediate queries.
In practice, visibility may be further reduced by Content Delivery Networks (CDNs), reverse proxies, or load balancers that aggregate multiple services under a small number of endpoints. To capture such weaker settings, we explicitly evaluate partial-visibility scenarios in which only a fraction of domains (e.g., 80%–10%) is observable (Table 11). These capabilities correspond to realistic on-path observers such as local ISPs, corporate firewalls, DNS resolvers, VPN providers, or shared network operators that collect encrypted metadata without decryption or manipulation.
Adversary Assumptions. We assume the user interacts with a local web or research agent that autonomously performs web browsing in response to a natural language prompt. The adversary is aware of the agent’s general browsing capabilities, such as issuing web searches, clicking links, scrolling pages, or scraping structured data from web pages. The adversary monitors the session between the agent and the Internet but does not alter traffic. We treat full-session metadata visibility as an upper bound: weaker adversaries such as DNS resolvers or enterprise firewalls may see only subsets of domains, yet our experiments show that WRA traffic remains highly distinguishable due to its structured, high-volume access patterns. We further assume that DNS and TLS handshake metadata remain visible (i.e., no DoH, DoT, or ECH) and that users do not employ VPNs, Tor, or other anonymizing proxies that would conceal domain sequences completely.
@@ 3.2 Attack Scenarios
Academic WRAs. In university, students and faculty use local WRAs for daily tasks, coursework, or research, leaking information about politics or unpublished work. External observers like ISPs or competing research entities could exploit traces to reconstruct prompts or profile academic interests.
@@ 4 Attack Methodology
We present the end-to-end pipeline for our privacy attacks against AI agents. Both tasks—prompt recovery and trait inference—follow a shared two-stage process: (1) offline model construction on proxy datasets or , yielding or ; and (2) online attack execution on real user traces or to produce or . The pipeline remains structurally consistent across tasks, though the input sources and inference targets differ.
@@ 4.1 Proxy Prompt Generation
To train the inference model in Stage 1, the adversary generates proxy prompts that realistically simulate user behavior and produce informative traces . Effective proxy prompts must balance realism, browsing induction, and inferential utility. (a) Prompts resemble natural-language queries real users might issue (health, education, planning). (b) They must reliably trigger multi-page browsing, rather than being satisfied by a single source (e.g., Wikipedia) or by the agent’s prior knowledge; to enforce this, we prepend instructions requiring at least five distinct page visits. (c) Prompts must be sufficiently detailed to produce distinguishable traces; underspecified inputs are rewritten into instruction-complete forms to avoid trivial queries that yield little signal. (d) Prompts and resulting traces remain semantically linked, so the trace encodes sufficient signal for inference.
Prompt recovery. We draw proxy prompts from public query datasets (FEDWEB13 [fedweb13], SESSION14 [trec14-session-data], DD16 [trec16-domain-data]) to build . Some prompts are underspecified to reliably induce meaningful browsing, and agents differ in how they handle such inputs. GPT-Researcher [elovic2023gptresearcher] internally refines vague prompts, but ChatGPT Deep Research UI asks clarifying questions, while Deep Research API, AutoGen, and Browser-Use execute inputs as-is. To align behaviors, Deep Research API prompts are rewritten into detailed variants (denoted -DR) to match the interactive interface.
Trait inference. From persona profiles, we construct , embedding selected traits either explicitly or implicitly into prompts. This mirrors real user behavior, where personal context guides information seeking (e.g., health questions tied to age or history). A persistent adversary linking sessions over time (e.g., by IP) would observe recurring patterns tied to these traits (). Executing such prompts through agents produces labeled traces that reveal stable user attributes, enabling training of trait inference models (). Dataset-specific sampling and generation details are described in Section 6.1.
@@ 4.2 Domain Trace Collection
To characterize how different prompts induce distinct browsing behavior, we collect domain-level metadata from agents via Playwright or built-in system logs. For each prompt , this yields a trace , where denotes accessed domain, timestamp, and payload size. The ordered trace captures whether an agent revisits the same domains or explores the set of sources during a single prompt execution. Although TLS application data does not expose full URLs or domain names directly, domain attribution remains feasible through DNS queries, SNI fields in TLS handshakes, or correlated DNS/HTTPS timing, all of which are observable to ISPs or other on-path entities. This matches realistic adversarial capabilities today, though we note that CDN consolidation or future deployment of DoH/DoT/ECH could reduce visibility.
Noise Filtering. A passive adversary observes all outbound requests, including primary domains, embedded resources, ads, and trackers. Because much of this traffic is incidental rather than agent-driven, an effective adversary would attempt to filter noise to isolate the domains most indicative of user intent. We apply domain-based noise filtering to remove well-known advertising and tracking services (e.g., doubleclick.net, onetrust.com, taboola.com). This represents the simplest plausible filtering approach and can be implemented using standard blocklists (e.g., EasyList [easylist2025]). More advanced adversaries could employ refined techniques (e.g., traffic classification or resource analysis) to further separate incidental requests from meaningful domains, which we leave to future work.
@@ 4.3 Inference Model Construction and Use
For prompt recovery, the task is to reconstruct a natural-language prompt that captures the functional intent of the user’s original input. The adversary constructs an inference model () from proxy (trace, prompt) pairs in the proxy dataset (), using either ICL, where a small set of examples is included in the inference prompt without parameter updates, or fine-tuning, where the same proxy pairs are used to train a direct mapping from traces to prompts (Table 2). Given a user prompt producing trace (optionally filtered to per Section 4.2), the resulting model is then applied to new traces to reconstruct real user prompts .
@@
5 Ontology-aware Behavioral Leakage Score
We introduce OBELS (Ontology-aware Behavioral Leakage Scores), a multi-metric scoring scheme that decomposes leakage into dimensions of preserved functionality, domain targets, and semantic entities, providing a structured alignment with attacker goals and real-world agent behavior.
@@ 5.1 Limitations of Existing Metrics
In AI agent contexts, leakage risk stems from functional equivalence: prompt recovery is risky if it leads the agent to similar sites, services, or goals, despite differences in wording. Existing metrics fall short because they assess what prompts say, not what they do.
SBERT score, computed as cosine similarity between sentence embeddings [reimers2019sentence], captures high-level semantic alignment but suffers from paraphrastic inflation. It misleadingly assigns high scores to verbose, generic, or loosely related prompts [steck2024cosine, you2025semantics], overlooking task-specific behavior or intent.
LLM-as-Judge methods leverage LLMs to rate similarity through natural language instructions. While more tolerant to paraphrase and flexible, they are prompt-sensitive and often lack interpretability [razavi2025benchmarking, lee2024one], producing scalar scores without indicating which semantic elements were preserved or lost.
These shortcomings are evident in practice. For example, rephrasing “Compare Italy and Spain’s digital nomad visas” as “How do work visas differ across southern Europe?” yields low cosine similarity, yet the agent navigates to similar government portals and retrieves overlapping content. From a privacy standpoint, this is a successful reconstruction, yet traditional metrics fail to capture it.
@@ 5.2 Semantic Triplet Scoring with OBELS
To address these gaps, we propose OBELS, a set of four complementary scores designed to capture behavioral alignment. Rather than collapsing leakage into a single scalar, OBELS decomposes alignment into distinct dimensions so that failures in one area (e.g., entity granularity) are not masked by strong similarity elsewhere. This decomposition yields interpretable diagnostics and aligns more directly with attacker goals and real-world agent behavior. It does so by abstracting prompts into structured triplets:
@@
Here, intent denotes the high-level goal (e.g., learn, explore, analyze, compare, summarize, plan, decide, book, watch, read, evaluate); source_type identifies the class of information or service sought (e.g., travel, symptom, policy_area, event, visa_process, treatment_method, academic_field, cuisine, etc); and entity specifies the core topic (e.g., Rome, Italy, depression, face transplant, immigration, PhD, Business, Swahili dish, etc). Removing any triplet component collapses distinct browsing behaviors, whereas variations in tone or sentiment do not meaningfully alter domain traversal.
@@
Each dimension is scored from 0.0 (unrelated) to 1.0 (fully equivalent) using a standardized evaluation template applied to GPT-4, which compares triplet sets holistically and assigns fine-grained similarity scores.
Different applications may weight these components differently. For instance, entity granularity and semantic precision dominate in high-risk healthcare scenarios; domain-type alignment matters most in regulatory tasks; and intent alignment is sufficient for lower-risk consumer scenarios such as travel planning. Collapsing these dimensions into a single score would obscure such task-sensitive variation and misrepresent leakage risk.
@@
| FEDWEB13 [trec13-fedweb-topics] | Prompt | 50 prompts | 20 | GPT-Researcher | 77.5 | 143.0 | 1 prompt 1 session |
| SESSION14 [trec14-session-data] | Prompt | 60 prompts | 20 | GPT-Researcher | 69.9 | 145.9 | 1 prompt 1 session |
| DD16 [trec16-domain-data] | Prompt | 53 prompts | 20 | GPT-Researcher | 64.8 | 155.8 | 1 prompt 1 session |
| PERSONA [castricato2024persona] | Trait | 50 personas | 3 | AutoGen | 144.1 | 173.7 | 3–5 prompts 7 sessions |
@@ 6.1 Datasets
Prompt Recovery (TREC). We use three TREC datasets: FedWeb 2013 (50 topics) [trec13-fedweb-topics], Session 2014 (60) [trec14-session-data], and Dynamic Domain 2016 (53) [trec16-domain-data], referred to as FEDWEB13, SESSION14, and DD16. For FEDWEB13 and DD16, we concatenate the <description> and <narrative> fields to match SESSION14’s instruction-style and prompt specificity. To align with OpenAI’s Deep Research API [openai2025introducingdeepresearch], we also generate -DR variants by rewriting each prompt with GPT-4.1, while scoring is always against the original prompts. For each dataset, 20 prompts are reserved for testing, with the remainder used for training and ICL examples.
TREC query text may appear in LLM pre-training corpora, but the domain-level traces we evaluate on are newly collected and do not exist in any pre-training dataset. The attack exploits these fresh behavioral traces, not memorized prompt content.
Trait Inference (PERSONA). We sample 50 profiles from PERSONA_subset [castricato2024persona] (997 total), each annotated with 32 traits. For each, we select 5 traits and use GPT-4o to generate 3–5 prompts across 7 sessions, simulating a week of browsing. Three personas are held out for ICL, and after filtering to ages 18–70, 35 remain for testing.
@@ 6.2 WRAs and Backbone LLMs
GPT Researcher [elovic2023gptresearcher]. An open-source autonomous research agent for deep multi-hop investigation. We run it in deep mode, which produces structured research traces. GPT Researcher strategically uses multiple LLM backbones, and we evaluate both GPT-family and open-source backbones for the prompt recovery task.
Browser-Use [browser_use2024]. An open-source web agent that automates browsing through a local browser. To ensure stable traces, we configure it to start from Bing and require at least five distinct page visits before summarization. Browser-Use is used for the prompt recovery task with GPT-4o as the backbone.
AutoGen [wu2023autogen]. An open-source web agent system we used for both prompt recovery (GPT-4o) and trait inference (Gemini 2.0 Flash), instrumented to log domains and metadata. Unlike the other agents, it produces the real packet-level traces, noisier with ads and trackers.
@@ 6.3 Evaluation Metrics
(1) SBERT, computed as cosine similarity over mean-pooled Sentence-BERT embeddings [reimers2019sentence];
(2) LLM-as-Judge
, where an LLM rates prompt similarity based on natural language instructions; and
(3) OBELS, our proposed behavior-aware scoring scheme to capture functional equivalence (i.e., whether the reconstructed prompt would induce similar agent behavior).
@@
Categorical traits use exact match for single-token values; multi-word categories are scored using SBERT to account for semantic equivalence.
Free-text traits are assessed using SBERT, which computes semantic overlap via contextual token embeddings.